No need for excel handwritten answers will be enoughEXCEL Question 3 (Annuity with decreasing payments) (3marks] Construct an EXCEL spreadsheet, which projects cash flows of an
Answer #1
Year Cash flow
1 15000
2 14750
3 14500
4 14250
5 14000
6 13750
7 13500
8 13250
9 13000
10 12750
11 12500
12 12250
13 12000
14 11750
15 11500
16 11250
17 11000
18 10750
19 10500
20 10250
21 10000
22 9750
23 9500
24 9250
25 9000
Present value $                   1,62,757.05

B138 fax А Year 111 112 2 3 113 114 115 116 117 118 119 120 121 123 Cash flow 15000 =B112-250 =B113-250 =B114-250 =B115-250 =